OpenStudy (anonymous):

Can the set of lengths be the side lengths of a right triangle? 7ft, 12 ft, 17ft

OpenStudy (alexandervonhumboldt2):

check using pythgorian theorem: a^2+b^2=c^2 7^2+12^2=17^2 49+144=289 193=289 which is false, thus cannot be

OpenStudy (anonymous):

thank you